GEODESIC SURFACES IN KNOT COMPLEMENTS

IAIN R. AITCHISON AND J. HYAM RUBINSTEIN Abstract. We construct closed, embedded, totally geodesic surfaces in two hyperbolic manifolds obtained as knot complements.